Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability doubt.

Expert replies
by Aishwarya1204 » Sun Oct 14, 2012 1:16 am
Hey can somebody please help me with this question.
I faced it while giving the Kaplan practice test.

Q: The events A and B are independent. The probability that event A occurs is p and the probability that event B occurs is q. Which of the following is equal to the probability that exactly one of the events A or B occur.

p-2pq
p-pq
p+q-pq
p+q-2pq
p+q


Now, what I did was use the formula : p(that event A occurs) + p(that event B occurs) - p(both events occur)
which gave p + q - pq

But the answer is actually p + q - 2pq.

Thanks.
Join the discussion
Source: — Problem Solving |

by Anurag@Gurome » Sun Oct 14, 2012 4:50 am
Aishwarya1204 wrote:Hey can somebody please help me with this question.
I faced it while giving the Kaplan practice test.

Q: The events A and B are independent. The probability that event A occurs is p and the probability that event B occurs is q. Which of the following is equal to the probability that exactly one of the events A or B occur.

p-2pq
p-pq
p+q-pq
p+q-2pq
p+q


Now, what I did was use the formula : p(that event A occurs) + p(that event B occurs) - p(both events occur)
which gave p + q - pq

But the answer is actually p + q - 2pq.

Thanks.
Probability that exactly one of the events A and B occurs means if A occurs then B does not occur or if B occurs then A does not occur.

Therefore, p * (1 - q) + q * (1 - p) = p - pq + q - pq
= p + q - 2pq

The correct answer is D.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by Aishwarya1204 » Sun Oct 14, 2012 5:16 am
Anurag@Gurome wrote:
Aishwarya1204 wrote:Hey can somebody please help me with this question.
I faced it while giving the Kaplan practice test.

Q: The events A and B are independent. The probability that event A occurs is p and the probability that event B occurs is q. Which of the following is equal to the probability that exactly one of the events A or B occur.

p-2pq
p-pq
p+q-pq
p+q-2pq
p+q


Now, what I did was use the formula : p(that event A occurs) + p(that event B occurs) - p(both events occur)
which gave p + q - pq

But the answer is actually p + q - 2pq.

Thanks.
Probability that exactly one of the events A and B occurs means if A occurs then B does not occur or if B occurs then A does not occur.

Therefore, p * (1 - q) + q * (1 - p) = p - pq + q - pq
= p + q - 2pq

The correct answer is D.

Thank You so much !
But can you tell me what the main difference is between the two. The formula ideally means that exactly one event occurs as well right?
Join the discussion

by Anurag@Gurome » Sun Oct 14, 2012 6:21 am
Aishwarya1204 wrote: Thank You so much !
But can you tell me what the main difference is between the two. The formula ideally means that exactly one event occurs as well right?
When we count the odds of A, and independently counting the odds of A means any time A and B both happen, which means we are counting two times.
Therefore, the odds of at least one of A and B occurring is p + q - pq - subtracting the overlap.

In this question, we subtract the overlap, pq, again. I hope it helps.
Anurag Mairal, Ph.D., MBA
GMAT Expert, Admissions and Career Guidance
Gurome, Inc.
1-800-566-4043 (USA)

Join Our Facebook Groups
GMAT with Gurome
https://www.facebook.com/groups/272466352793633/
Admissions with Gurome
https://www.facebook.com/groups/461459690536574/
Career Advising with Gurome
https://www.facebook.com/groups/360435787349781/
Join the discussion

by Brent@GMATPrepNow » Sun Oct 14, 2012 7:14 am
Aishwarya1204 wrote: Q: The events A and B are independent. The probability that event A occurs is p and the probability that event B occurs is q. Which of the following is equal to the probability that exactly one of the events A or B occur.
p-2pq
p-pq
p+q-pq
p+q-2pq
p+q
If events A and B are independent, then P(A and B) = P(A) x P(B)

Now let's reword the probability. To do this, I always ask, "What needs to occur here to achieve the stated outcome?"

We get: P(A only OR B only) = P(A and not B _ OR _B and not A)
= P(A and not B) + P(B and not A)
= [P(A) x P(not B)] + [P(B) x P(not A)]

Aside: If P(A) = p, then P(not A) = 1 - p
Similarly, P(B) = q, then P(not B) = 1 - q

Now replace the given probabilities to get:
= [p x (1-q)] + [q x (1-p)]
Simplify to get: [p - pq] + [q - pq]
= p + q - 2pq
= D

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion